# Ditch the Denoiser: Emergence of Noise Robustness in Self-Supervised Learning from Data Curriculum

### v1

@@ B.3 Restart at Different Epochs
We investigated the impact of restarting training on noisy data at different stages: (1) training on denoised data for 140 epochs, followed by 60 epochs on noisy data (a common setting in this paper); (2) 130 epochs on denoised data, then 70 epochs on noisy data; and (3) 120 epochs on denoised data, then 80 epochs on noisy data. As shown in Figure 11, the performances across these variations are similar, with no approach consistently outperforming the others. Thus, we decide to use 140 epochs as the restart point.
@@ B.4 Enduring Effectiveness in Mixed Noisy-Clean Data
In real-world scenarios, datasets often contain images of varying quality, posing challenges for SSL methods. In this section, we investigate how different noise-clean distributions on ImageNet-100 affect the SSL model’s performance and the effectiveness of our method under these distributions. We replace $x\%$ (i.e., $0\%,2\%,10\%$) of noisy data with clean images. So in our noise curriculum, the model is trained on a mix of $(100-x)\%$ denoised and $x\%$ clean images in the first stage, then on a mix of $(100-x)\%$ noisy and $x\%$ clean images in the second stage. The validation set is also adjusted to follow the same distribution as the training set. From Table 8, we can see that generally the linear evaluation accuracy decreases across all methods when more clean data is introduced but noisy data remains the majority, which can be attributed to a reduced overall representation alignment between noisy, denoised and actual clean data. On the other hand, through column-wise comparisons, DINOv2 w/ NC consistently outperforms DINOv2 and closely matches N2N + DINOv2; regularized method DINOv2 w/ NCT brings substantial improvement over unregularized method. It is clear that our curriculum learning framework and denoised regularization are still effective despite the varying training distributions.
@@
Table 8: Mean Average Precision (mAP) on instance-level recognition tasks for other SSL models. The pattern is consistent with Table 3. Noise curriculum (NC) improves performance for all models.
@@ B.5 Further Results for Other SSL Models
Figure 12 shows the accuracy over epochs for SSL models tested in Section 4.6. Table 8 shows the performance of the SSL models on the instance recognition task using the Oxford and Paris dataset. Both results illustrate that the noise curriculum improves the performance of all SSL models on noisy data, demonstrating the wide applicability of the method.
@@
Appendix C Visualizations
C.1 Training Loss Visualization
@@
Figure 13 shows the comparison between training loss of DINOv2 and DINOv2 w/ NC on the ImageNet-100 noisy data with Gaussian noise ($\sigma=100$). DINOv2 w/ NC loss (green) descends to a lower value than DINOv2 (red). The two curves are similar in shape which reflects the restarting training technique in our noise curriculum learning.
C.2 Noisy Images Visualization

### v2

@@ B.3 Restart at Different Epochs
We investigated the impact of restarting training on noisy data at different stages: (1) training on denoised data for 140 epochs, followed by 60 epochs on noisy data (a common setting in this paper); (2) 130 epochs on denoised data, then 70 epochs on noisy data; and (3) 120 epochs on denoised data, then 80 epochs on noisy data. As shown in Figure˜11, the performances across these variations are similar, with no approach consistently outperforming the others. Thus, we decide to use 140 epochs as the restart point.
@@ B.4 Enduring Effectiveness in Mixed Noisy-Clean Data
In real-world scenarios, datasets often contain images of varying quality, posing challenges for SSL methods. In this section, we investigate how different noise-clean distributions on ImageNet-100 affect the SSL model’s performance and the effectiveness of our method under these distributions. We replace $x\%$ (i.e., $0\%,2\%,10\%$) of noisy data with clean images. So in our noise curriculum, the model is trained on a mix of $(100-x)\%$ denoised and $x\%$ clean images in the first stage, then on a mix of $(100-x)\%$ noisy and $x\%$ clean images in the second stage. The validation set is also adjusted to follow the same distribution as the training set. From Table˜8, we can see that generally the linear evaluation accuracy decreases across all methods when more clean data is introduced but noisy data remains the majority, which can be attributed to a reduced overall representation alignment between noisy, denoised and actual clean data. On the other hand, through column-wise comparisons, DINOv2 w/ NC consistently outperforms DINOv2 and closely matches N2N + DINOv2; regularized method DINOv2 w/ NCT brings substantial improvement over unregularized method. It is clear that our curriculum learning framework and denoised regularization are still effective despite the varying training distributions.
@@
Table 8: Mean Average Precision (mAP) on instance-level recognition tasks for other SSL models. The pattern is consistent with Table˜3. Noise curriculum (NC) improves performance for all models.
@@ B.5 Further Results for Other SSL Models
Figure˜12 shows the accuracy over epochs for SSL models tested in Section˜4.6. Table˜8 shows the performance of the SSL models on the instance recognition task using the Oxford and Paris dataset. Both results illustrate that the noise curriculum improves the performance of all SSL models on noisy data, demonstrating the wide applicability of the method.
@@
B.6 Robustness to Denoiser Quality
To investigate the level of tolerance of DINOv2 w/ NC to the denoiser quality, we conduct ablation studies on how the number of denoiser training epochs impacts the SSL model’s final performance. We use the number of training epochs as a quantifiable proxy for denoiser effectiveness. The evaluation follows our default setting: 200 epochs of DINOv2 ViT-S training with Gaussian-100 noisy data. The results are shown in Table˜9. We observe that even training the N2N denoiser for just 1 epoch leads to a substantial improvement of DINOv2 w/ NC over the DINOv2 baseline (i.e., 10.7 increase), while being only 2 percent less than that of a well-trained denoiser (100 epochs). Although the 1-epoch denoiser still produces many undesirable artifacts like a very weak denoiser, the curriculum still yields substantial improvements in downstream performance. This highlights the robustness of our curriculum, which has considerable tolerance to the denoiser’s quality. A denoiser has to be unreasonably bad (e.g., training diverged or collapsed) to provide no return in performance.
Table 9: Performance of DINOv2 w/ NC on ImageNet-100 with the N2N denoiser trained for varying epochs. Notably, even a weak denoiser trained for only 1 epoch substantially improves DINOv2 w/ NC over DINOv2.
Method Denoiser Training Epochs Accuracy
DINOv2 / 55.4
DINOv2 w/ NC 1 epoch 66.1
DINOv2 w/ NC 5 epochs 67.4
DINOv2 w/ NC 100 epochs 68.1
Appendix C Baseline Comparisons
C.1 Performance of Clean-Pretrained Model on Noisy Data
We evaluate the noise-robustness of a DINOv2 model trained exclusively on clean images, in order to reinforce the necessity of noisy training. Two strategies are employed to cover complementary baselines. (1) With the backbone frozen, we fine-tune a linear probe on noisy data drawn from the same distribution as the test set, and then evaluate the full model on noisy data. (2) With the backbone again frozen, we fine-tune a linear probe on clean data and subsequently evaluate the model on noisy data. The first strategy simulates a common downstream scenario where a frozen feature extractor is adapted to a new domain via task-specific heads, while the second serves as a more standard adversarial benchmark that tests the model’s robustness to noise without ever exposing it to noise during training. In addition to our own trained models, we also benchmark the officially released DINOv2 weights, which were pretrained on the LVD-142M dataset (142 million images). The results in Table˜10 show substantial performance drops (15–70 points) when applying clean-pretrained DINOv2 to noisy test sets. The degradation is particularly severe on Gaussian-255 noise, where the accuracy of ViT-B/14 drops from 84.5 to 29.0 (a 55.5-point decrease), even when the probe is tuned on Gaussian-255 noisy data.
Notably, despite the smaller model size and significantly reduced training data, our DINOv2 w/ NCT (Table 1(b)) surpasses the official ViT-B/14 weights trained on LVD-142M at both Gaussian-100 (72.1 vs. 70.3) and Gaussian-255 (55.8 vs. 29.0). These results emphasize that standard DINOv2 is not robust to input noise, underscoring the necessity of incorporating noisy images during pretraining to achieve strong downstream performance in noisy settings.
Table 10: Performance of clean-pretrained DINOv2 models under different noise conditions. Linear probes are fine-tuned either on clean or noisy data while keeping the backbone frozen. The clean-pretrained DINOv2 suffers severe drop in performance when applied on the noisy test set.
Training Dataset Architecture Eval Dataset Eval Noise
Probe Tuned
on Clean
Probe Tuned
on Noisy
LVD-142M ViT-B/14 ImageNet-1k Gaussian-100 63.2 70.3
Gaussian-255 8.9 29.0
Clean 84.5 /
LVD-142M ViT-S/14 ImageNet-1k Gaussian-100 49.2 61.0
Gaussian-255 4.9 23.1
Clean 81.1 /
ImageNet-1k
(100 epochs)
ViT-B/16 ImageNet-1k Gaussian-100 40.7 58.1
Gaussian-255 2.2 20.5
Clean 79.0 /
ImageNet-100
(1000 epochs)
ViT-S/16 ImageNet-100 Gaussian-50 60.0 74.3
Gaussian-100 22.3 61.3
Gaussian-255 3.1 34.5
Clean 81.4 /
C.2 Baseline Performance on Mixed Noisy–denoised Data
Table 11: Linear probing accuracy on noisy ImageNet-100 test set when training on mixed noisy+denoised data. Mixing denoised images results in a 7.4 point drop compared to training on noisy data alone.
Training Setting Accuracy
100% Gaussian-100 55.4
50% Gaussian-100 + 50% Denoised 48.0
Another intuitive baseline is to investigate whether mixing denoised samples with noisy ones during training can improve representation learning by increasing the diversity and expressiveness of inputs. Table˜11 shows the linear probing accuracies of DINOv2 ViT-S on the noisy ImageNet-100 test set, when trained for 200 epochs on 100% noisy data (from Table 1(a)) versus on 50% noisy + 50% denoised data. We observe a significant drop of 7.4 points in accuracy when noisy and denoised data are randomly mixed during training. As a result, mixing clean (denoised) and noisy images during training leads to degraded performance, primarily due to reduced representation alignment between noisy and denoised samples, which hinders the model’s ability to learn consistent features. This shows naive mixing is suboptimal, and reinforces the necessity of using a staged curriculum learning approach that separates noisy and denoised images.
Appendix D Visualizations
D.1 Training Loss Visualization
@@
Figure˜13 shows the comparison between training loss of DINOv2 and DINOv2 w/ NC on the ImageNet-100 noisy data with Gaussian noise ($\sigma=100$). DINOv2 w/ NC loss (green) descends to a lower value than DINOv2 (red). The two curves are similar in shape which reflects the restarting training technique in our noise curriculum learning.
D.2 Noisy Images Visualization
@@
D.3 PCA Visualization
To visualize the dense features, we apply PCA to the feature space and map the first three principal components to the RGB channels. We use the ViT-S 200-epoch Gaussian-255 checkpoints from Table˜1(a) for visualization. As shown in , the DINOv2 visualizations (middle of each triplet) fail to clearly separate salient objects from the background, and the resulting color patterns lack semantic consistency across images, suggesting weaker representation quality. In contrast, the DINOv2 w/ NCT visualizations (right of each triplet) distinctly separate salient objects, and objects of the same class are consistently represented with similar color patterns across different images, indicating semantically aligned features. These results demonstrate that NCT yields a substantial improvement in feature quality over the DINOv2 baseline.
Figure 23: PCA visualizations: we visualize dense features of the models by performing PCA on the feature space and mapping the top three principal components to RGB. In each triplet, the left image is the noisy input with Gaussian noise at $\sigma=255$. The middle and right images are produced by DINOv2 and DINOv2 w/ NCT (ViT-S, 200 epochs) respectively.